Cos47÷sin43+sin72÷cos18-2cos square 45 evaluate the value

Grade:9

1 Answers

Shailendra Kumar Sharma
6 years ago
Cos47/sin (90-47)+Sin72/cos(90-72) -2(1/2)
Since Cos (90-A) =SinA and Sin (90-A) =CosA
1+1-1 =1 Answer